OpenStudy (anonymous):

Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? Stepsss!!!

OpenStudy (anonymous):

To find the other two sides when all you know is one angle and one side you compare it to the various trig functions. 

OpenStudy (anonymous):

These are just the definitions for the trig functions: sin b = cos a = q/h cos b = sin a = p/h tan a = p/q tan b = q/p You just look for the equation that has the variable you're trying to find along with two that you know. So if we know a is 30 degrees, and we know h is 10. We can find side q. cos 30 = q/10 10 cos 30 = q or we could find side p sin 30 = p/10 10 sin 30 = p. And so on from there.

OpenStudy (anonymous):

Okay thanksss